# KAN-ODEs: Kolmogorov-Arnold Network Ordinary Differential Equations for Learning Dynamical Systems and Hidden Physics

### v2

@@ III.2 Modeling Hidden Physics in PDEs: Fisher-KPP PDE
Note that this problem setup was adopted from [9]. The spatial grid was discretized with $\Delta x=0.04$, and the equation was solved with a time step of $\Delta t=0.5$. The obtained solution field $u(x,t)$ is shown in Fig. 5(A) and used as training data.
@@ III.3.1 Burgers’ Equation
In this formulation, the KAN works as a non-linear operator for the partial differential equation (PDE) solution. The KAN-ODE was constructed with 2 layers: [51, 10, 5], [10, 51, 5]. Ten nodes in the intermediate layer were used to preserve the high-dimensional information. Note that the numbers of the input and output dimensions of the KAN were determined by the number of discretized state variables in $\mathbf{u}$. The KAN parameters were updated with a learning rate of 0.01 with the ADAM optimizer.
@@ III.3.2 Schrödinger Equation
where $u\in\mathbb{C}^{1}$. The time and spatial domains were set to $x\in[-5,5]$ and $t\in[0,\pi/2]$. The initial condition and boundary conditions were defined as following in Eqs. 24-25. Note that the problem settings used in this example were identical with those in [6].
@@ IV Conclusions
All the associated codes accompanying this manuscript will be made publicly available upon the acceptance of the manuscript.
@@ Acknowledgement
The work is supported by the National Science Foundation (NSF) under Grant No. CBET-2143625. BCK is partially supported by the NSF Graduate Research Fellowship under Grant No. 1745302.
@@ Appendix A. Sparse Regression on Lotka-Volterra KAN-ODE
Symbolic functions corresponding to the Symbolic KAN-ODE plotted in Fig. 4(B) are provided in Table A1. See Fig. 4(B) for a visual representation of the twelve activations and of the hidden layer nodes $n_{1}$ through $n_{3}$. The selection of candidate functions was carried out to balance accuracy with sparsity. For example, a linear term suffices to capture the linear behavior of $l1$ connecting $x$ and $n_{1}$, while $l2$ connecting $x$ and $n_{2}$ requires an additional $1/x$ term due to its strong nonlinear behavior. Similarly, the relatively complex $l11$ connecting $n_{3}$ to $dx/dt$ uses a linear term in addition to a polynomial division term. All activations were derived independently based on the input/output pairs present in the training data.
Table A1: Symbolic KAN for Lotka-Volterra dynamics. The first layer takes the input of $x$ and $y$ while the second layer takes the input of $n_{1}$, $n_{2}$, and $n_{3}$. Note that $n_{i}(x,y)=n_{i}(x)+n_{i}(y)$ and $(dx/dt)(n_{1},n_{2},n_{3})=(dx/dt)(n_{1})+(dx/dt)(n_{2})+(dx/dt)(n_{3})$ by definition in the KAN.
@@ Appendix B. Varying KAN-ODE Architectures for Fisher-KPP and Burgers’ Equations
Additional tests were performed to provide KAN-ODE performance metrics for varying KAN architectures in the Fisher-KPP equation and Burgers’ equation examples. Table B1 shows the test matrix with different sizes of KANs. For the Fisher-KPP equation, a relatively small KAN with very few parameters was used to model the hidden physics in the interpretable form (i.e. replacing only the source term), while larger KANs with more parameters (thanks to the larger input/output dimension) were constructed for the Burgers’ equation to build a surrogate model that predicted the entire solution field. Figure B1 illustrates the training losses with respect to the number of parameters in these two examples. We surprisingly found in both cases that small architectures of KANs, i.e. those with zero to one hidden node and three to five grid points, can successfully learn these dynamical systems with a training loss smaller than 10-6. Furthermore, in most cases, KAN-ODEs have a scaling rate of loss $\mathcal{L}\propto N^{-4}$, where $N$ is the total number of parameters in the KAN. This result empirically supports the validity of the scaling law in the original KAN [26], as well as results seen in the Lotka-Volterra example shown in Fig. 3(C).
Table B1: KAN-ODEs architecture tested for performance evaluation with the Fisher-KPP equation and the Burgers’ equation.
@@ C.1. Modeling Hidden Physics in PDEs
Figure C1: Allen-Cahn equation: (A) Solution field $u(x,t)$ of ground truth. (B) Solution field $u(x,t)$ of prediction with a learned KAN-ODE model. (C) Loss function. (D) Learned hidden physics of the reaction source term in the Allen-Cahn equation and its symbolic expression.
The KAN-ODE model was trained on the data shown in Fig. C1(A). In this example, we applied an early stop when the error reached 10-7. The prediction with the learned model is illustrated in Fig. C1(B), showing accurate reconstruction of the solution field $u(x,t)$ after $\sim$65,000 updates (Fig. C1(C)). This resultant single activation function is shown in Fig. C1(D). For the modeled activation function, a symbolic expression was derived, using the SymbolicRegression.jl package [42]. Given the candidate symbolic expressions of [+,,×,/][+,-$\times$,/][ + , - , × , / ], it returned the expressions tabulated in Table. C1. Among them, the following expression showed good accuracy,
@@ C.2. Data-Driven Solutions of PDEs
We further demonstrate KAN-ODEs as surrogate models using the Allen-Cahn equation by replacing the entire right-hand side of Eq. C1 with a KAN. To mimic sparse data measurements, only five profiles at the selected times $t\in\{0.1,0.3,0.5,0.7,0.9\}$ were collected from the ground truth simulation and used as training data (Fig. 22(A)), while testing involved reconstruction of the entire time history. The KAN-ODE as a surrogate model was constructed using 2 layers with 10 hidden nodes ([41,10,10], [10,41,10]). Equation 22 was trained to learn the dynamics of the phase transition as shown in the loss plot (Fig. C2(B)). The resulting field reconstruction of $u(x,t)$ is illustrated in Fig. C2(C), which qualitatively matches well with the ground truth. The spatio-temporal error between the KAN-ODE model and the ground truth ($u^{\text{KAN}}(x,t)-u^{\text{true}}(x,t)$) in Fig. C2(D) depicts near-zero error at and immediately near each training time slice, and small but noticeable error in the large windows between time steps, especially near the end of the simulation time. For example, the error is not appreciable at time $t$ = 0.1, but begins to grow before shrinking back to near-zero by $t$ = 0.3 when the second observation takes place. This result implies that the KAN-ODE needs more training data in order to extrapolate quantitatively well to times significantly outside of its training data window, like $t$ = 1.5 or $t$ = 2.0. It does, however, highlight the strong capability of KAN-ODEs to accurately model and reconstruct physical phenomena between training samples with very sparse measurements and without any prior physical knowledge included.
Figure C2: Allen-Cahn equation: Demonstration of KAN-ODEs as a surrogate model. (A) Ground truth $u(x,t)$. The dashed lines indicate the training data at $t$ = 0.1, 0.3, 0.5, 0.7, 0.9. (B) MSE loss during training. (C) Reconstructed $u(x,t)$ by the trained KAN-ODE model. (D) The error between the reconstructed $u^{\text{KAN}}(x,t)$ and the ground truth $u^{\text{true}}(x,t)$.

### v3

@@ III.2 Modeling Hidden Physics in PDEs: Fisher-KPP PDE
Note that this problem setup was adopted from [9]. The spatial grid was discretized with $\Delta x=0.04$, and the equation was solved with a time step of $\Delta t=0.5$ using the ODE integrator of Tsit5 (Tsitouras 5/4 Runge-Kutta method [41]). The obtained solution field $u(x,t)$ is shown in Fig. 5(A) and used as training data.
@@ III.3.1 Burgers’ Equation
In this formulation, the KAN works as a non-linear operator for the partial differential equation (PDE) solution. The KAN-ODE was constructed with 2 layers: [51, 10, 5], [10, 51, 5]. Ten nodes in the intermediate layer were used to preserve the high-dimensional information. Note that the number of input and output dimensions of the KAN were determined by the number of discretized state variables in $\mathbf{u}$. The KAN parameters were updated with a learning rate of 0.01 with the ADAM optimizer.
@@ III.3.2 Schrödinger Equation
where $u\in\mathbb{C}^{1}$. The time and spatial domains were set to $x\in[-5,5]$ and $t\in[0,\pi/2]$. The initial condition and boundary conditions were defined as following in Eqs. 24-25. Then, the governing equations were solved with a time step of $\Delta t=0.01$ and grid size of $\Delta x=0.05$ using the stiff ODE integrator of Rodas5 (A 5th order A-stable stiffly stable Rosenbrock method with a stiff-aware 4th order interpolant [49, 39]). Note that the problem settings used in this example were identical to those in [6].
@@ IV Conclusions
The KAN-ODE codes used to generate the results presented in this work are publicly available at https://github.com/DENG-MIT/KAN-ODEs.
@@ Acknowledgement
This work is supported by the National Science Foundation (NSF) under Grant No. CBET-2143625. BCK is partially supported by the NSF Graduate Research Fellowship under Grant No. 1745302.
@@ Appendix A. Sparse Regression on Lotka-Volterra KAN-ODE
Symbolic functions corresponding to the Symbolic KAN-ODE plotted in Fig. 4(B) are provided in Table A1. See Fig. 4(B) for a visual representation of the twelve activations and of the hidden layer nodes $n_{1}$ through $n_{3}$. The selection of candidate functions was carried out to balance accuracy with sparsity. For example, a linear term suffices to capture the linear behavior of $l1$ connecting $x$ and $n_{1}$, while $l2$ connecting $x$ and $n_{2}$ requires an additional $1/x$ term due to its strong nonlinear behavior. Similarly, the relatively complex $l11$ connecting $n_{3}$ to $dx/dt$ uses a linear term in addition to a polynomial division term. All activations were derived independently of each other, based on the immediately relevant univariate input/output pairs.
Table A1: Symbolic KAN activations for Lotka-Volterra dynamics. The first layer takes the inputs $x$ and $y$ while the second layer takes the inputs $n_{1}$, $n_{2}$, and $n_{3}$. Note that $n_{i}(x,y)=n_{i}(x)+n_{i}(y)$ and $dx/dt(n_{1},n_{2},n_{3})=dx/dt(n_{1})+dx/dt(n_{2})+dx/dt(n_{3})$ by definition in the KAN.
@@ Appendix B. Varying KAN-ODE Architectures for Fisher-KPP and Burgers’ Equations
Additional tests were performed to provide KAN-ODE performance metrics for varying KAN architectures in the Fisher-KPP equation and Burgers’ equation examples. Table B1 shows the test matrix with different sizes of KANs. For the Fisher-KPP equation, a relatively small KAN with very few parameters was used to model the hidden physics in the interpretable form (i.e. replacing only the source term). Larger KANs with more parameters (thanks to the larger input/output dimension) were constructed for the Burgers’ equation to build a surrogate model that predicted the entire solution field. Figure B1 illustrates the training losses with respect to the number of parameters in these two examples. We surprisingly found in both cases that small architectures of KANs, i.e. those with zero to one hidden node and three to five grid points, can successfully learn these dynamical systems with a training loss smaller than 10-6. Furthermore, in most cases, KAN-ODEs have a scaling rate of loss $\mathcal{L}\propto N^{-4}$, where $N$ is the total number of parameters in the KAN. This result empirically supports the validity of the scaling law in the original KAN [26], as well as results seen in the Lotka-Volterra example shown in Fig. 3(C).
Table B1: KAN-ODE architectures tested for performance evaluation with the Fisher-KPP equation and the Burgers’ equation.
@@ C.1. Modeling Hidden Physics in PDEs
Figure C1: Allen-Cahn equation: (A) Solution field $u(x,t)$ of the ground truth. (B) Solution field $u(x,t)$ of the prediction with a trained KAN-ODE model. (C) Loss function. (D) Learned hidden physics of the reaction source term in the Allen-Cahn equation, and its symbolic expression.
The KAN-ODE model was trained on the data shown in Fig. C1(A). In this example, we applied an early stop when the error reached 10-7. The prediction with the learned model is illustrated in Fig. C1(B), showing accurate reconstruction of the solution field $u(x,t)$ after $\sim$65,000 updates (Fig. C1(C)). This resultant single activation function is shown in Fig. C1(D). For the modeled activation function, a symbolic expression was derived, using the SymbolicRegression.jl package [45]. Given the candidate symbolic expressions of [+,,×,/][+,-$\times$,/][ + , - , × , / ], it returned the expressions tabulated in Table C1. Among them, the following was the simplest expression of the low-loss group (i.e. the group of expressions that jumped from $10^{-1}$ to $10^{-5}$ loss),
@@ C.2. Data-Driven Solutions of PDEs
We further demonstrate KAN-ODEs as surrogate models using the Allen-Cahn equation by replacing the entire right-hand side of Eq. C1 with a KAN. To mimic sparse data measurements, only five profiles at the selected times $t\in\{0.1,0.3,0.5,0.7,0.9\}$ were collected from the ground truth simulation and used as training data (Fig. C2(A)), while testing involved reconstruction of the entire time history. The KAN-ODE as a surrogate model was constructed using 2 layers with 10 hidden nodes ([41,10,10], [10,41,10]). Equation 22 was trained to learn the dynamics of the phase transition as shown in the loss plot (Fig. C2(B)). The resulting field reconstruction of $u(x,t)$ is illustrated in Fig. C2(C), which qualitatively matches well with the ground truth. The spatio-temporal error between the KAN-ODE model and the ground truth ($u^{\text{KAN}}(x,t)-u^{\text{true}}(x,t)$) in Fig. C2(D) depicts near-zero error at and surrounding each training time slice, and small but noticeable error in the large windows between time steps, especially near the end of the simulation time. For example, the error is not appreciable at time $t$ = 0.1, but begins to grow before shrinking back to near-zero by $t$ = 0.3 when the second observation takes place. While the errors introduced here are fairly minor, this result implies that the KAN-ODE needs more training data in order to extrapolate quantitatively well to times significantly outside of its training data window, like $t$ = 1.5 s or $t$ = 2.0 s. Within the 1 s range considered here, however, these results highlight the strong capability of KAN-ODEs to accurately model and reconstruct physical phenomena between training samples with very sparse measurements and without any prior physical knowledge included.
Figure C2: Allen-Cahn equation: Demonstration of a KAN-ODE as a surrogate model. (A) Ground truth $u(x,t)$. The dashed lines indicate the training data at $t$ = 0.1, 0.3, 0.5, 0.7, 0.9. (B) MSE loss during training. (C) Reconstructed $u(x,t)$ by the trained KAN-ODE model. (D) The error between the reconstructed $u^{\text{KAN}}(x,t)$ and the ground truth $u^{\text{true}}(x,t)$.